OpenStudy (anonymous):

how do you complete the square on this 7x^2 + 10x = -2

OpenStudy (amistre64):

by keeping your wits about you

OpenStudy (amistre64):

factor out the 7 on the left and work out the square part

OpenStudy (anonymous):

divide by 7 first?

OpenStudy (amistre64):

\[7x^2+10x=-2\] \[7(x^2+\frac{10}{7}x)=-2\] \[7(x^2+\frac{10}{7}x+(\frac{10}{14})^2)=-2+7(\frac{10}{14})^2\]

OpenStudy (amistre64):

then I spose we can get rid of that 7 by dividing both sides :)

OpenStudy (amistre64):

\[x^2+\frac{10}{7}x+(\frac{10}{14})^2=-\frac{2}{7}+(\frac{10}{14})^2\] compact the left \[(x+\frac{10}{14})^2=-\frac{2}{7}+(\frac{10}{14})^2\] and work the rest thru

OpenStudy (amistre64):

yes, divide off the 7 first :)

OpenStudy (anonymous):

i can get to the point where i get (x = (10/14))^2 = 44/196 but dont know what to do after this

OpenStudy (amistre64):

you sqrt the sides to unto the ^2 and add off the 10/14

OpenStudy (anonymous):

is the answer (-5+ Sqrt11)-14

OpenStudy (amistre64):

\[\sqrt{(x+\frac{10}{14})^2=-\frac{2}{7}+(\frac{10}{14})^2}\] \[x+\frac{10}{14}=\pm\sqrt{-\frac{2}{7}+(\frac{10}{14})^2}\] \[x=-\frac{10}{14}\pm\sqrt{-\frac{2}{7}+(\frac{10}{14})^2}\]

OpenStudy (anonymous):

i mean (-5 + sqrt11)/14

OpenStudy (amistre64):

the sqrt has 2 parts to it a + and a - that has to be accounted for

OpenStudy (amistre64):

x^2 = 4 x = sqrt(4) x = 2, or -2

OpenStudy (amistre64):

heres a simple dbl chk http://www.wolframalpha.com/input/?i=7x%5E2+%2B+10x+%3D+-2

OpenStudy (amistre64):

looks to be a /7 not /14

OpenStudy (anonymous):

ok thanks a mill

OpenStudy (amistre64):

yep